# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### v3

@@ 6.4 Performances in Cryogenic Storage’s Thermal Management
Figure 21: Performance of the digital twin assimilation. (a) Cost function $\mathcal{L}_{p}$ normalized with its initial value $\mathcal{L}_{p}(\bm{w}_{p}^{0})$ over the number of episodes. Agreement of the predicted state $\bm{s}=[V_{l},p,T_{s}]$ with the real state $\check{\bm{s}}$ shown through (b) a scatter plot, and (c) the time series evolution.
The agreement between the predicted state and the real one is presented in Figure 21, which shows both the time series of the real and predicted states (Figures 20(a) - 20(c)) and their direct comparison plotting one against the other (Figures 20(d) - 20(f)). All three components are successfully identified, even though slight deviations are found for $\bm{s}_{1}=p$ and $\bm{s}_{2}=T_{s}$.
The control performances of the optimal parametrization, $\bm{w}_{a}^{*}$, extracted from the learning curve presented in Figure 9(d) are studied next. The current scenario involves an initial over-pressurized tank operating 0.4 bar above the reference $\tilde{p}$. Figure 22 shows the time evolution of the real pressure $\check{p}$ and propellant mass $\check{m}_{l}$ over time, while the vented mass flow rate $\check{\dot{m}}_{c}$. Figures 21(b) and 21(c) show the control action over time and the corresponding pressure, respectively. Since the tank is initially in an over-pressurized state sufficiently far from $\tilde{p}$, the valve was regulated to open to its widest ($\dot{m}_{c}=0.63$ g/s), allowing the pressure to rapidly reduce until the target 5 bar are reached. This pressure drop is accompanied by a consumption of 87 kg over two roughly 45 hours. Once the target is reached, the active TVS regulates the pressure around this value. To do so, $\check{\dot{m}}_{c}$ fluctuates between -50% and 25% around its midpoint due to changes in the heat due to the disturbance, although such variations are not seen in the pressure signal.
Figure 22: Tracking performance of the pressure $\check{p}$, and variation of propellant mass $\check{m}_{l}$ over the real environment, starting from an over-pressure condition. Corresponding control action $\check{a}$ and vented mass flow rate $\check{\dot{m}}_{c}$. Thermodynamic actuation $\check{\dot{m_{v}}}$ and $\check{a}$ as a function of the pressure $p$. The acting policy $\pi(\check{\bm{e}};\bm{w}_{a})$ is retrieved by taking the optimal set of weights $\bm{w_{a}}$ from the training.
@@
Figure 23: Surface plot of the found policies, $\pi_{mf}$ (22(a)) and $\pi_{mb}$ (22(b)).
We inspect the two found policies in Figure 23. The two loops, model-based and model-free, converge to essentially the same parametrization. They show almost identical responses to the possible input space, with very minor discrepancies in the inclination of the hyperbolic tangent. Lastly, the robustness of the assimilation and control loops concludes this overview. The procedure consists in evaluating the cost of the optimal weights $\bm{w}_{p}$ and $\bm{w}_{a}$ over observations unseen during the training phase. To this end, we carried out one hundred simulations on random initial conditions and disturbances. For each of them, we computed the assimilation cost $\mathcal{L}_{p}$, as well as the control cost $\mathcal{L}_{a}$ using the model-based and model-free policies. This investigation, summarized in the violin plots presented in Figures 23(a) and 23(b), proves that the weights obtained at the end of the learning phase are valid and robust even in unseen scenarios. Moreover, in contrast to what was observed in the wind turbine test case, the model-based and model-free weights are nearly identical, thus their performances are also identical as supported by Table 2, which shows a difference between the two control performances in the order of $\mathcal{O}(10^{-2})$. This is likely a consequence of both approaches converging in the same region of the search space.
Figure 24: Violin plots obtained after 100 runs for the (a) assimilation performance and (b) model-based and model-free controllers.
@@ 7 Conclusions and Outlooks
The performance of the Reinforcement Twinning framework, both in training the digital twin and the controlling agent, proves the framework’s potential beyond any specific application. The merit stems from the complementary characteristics of the two approaches, the model-based being extremely sample-efficient but prone to fall stuck into local minima and the model-free being sample-inefficient but able to explore the parameter space better.
The paths towards future extensions are many. On the one hand, we could introduce a more advanced learning feedback mechanism between the model-free and the model-based methods, which is currently limited to a “cloning” procedure. This could allow, for example, different policy architecture between the two, unleashing the potential of reinforcement learning to handle arbitrarily complex policies in the form of an artificial neural network. On the other hand, we could introduce more sophisticated methods of integrating the digital twin in the training of the model-free loop, for instance, by using the virtual environment to boost the training of the critic network with strategies akin to traditional model-based reinforcement learning.

### v4

@@ 6.4 Performances in Cryogenic Storage’s Thermal Management
Figure 21: T3 - Performance of the digital twin assimilation. (a) Cost function $\mathcal{L}_{p}$ normalized with its initial value $\mathcal{L}_{p}(\bm{w}_{p}^{0})$ over the number of episodes. Agreement of the predicted state $\bm{s}=[V_{l},p,T_{s}]$ with the real state $\check{\bm{s}}$ shown through (b) a scatter plot, and (c) the time series evolution.
The agreement between the predicted state and the real one is presented in Figure 21, which shows both the time series of the real and predicted states (Figures 21(a) - 21(c)) and their direct comparison plotting one against the other (Figures 21(d) - 21(f)). All three components are successfully identified, even though slight deviations are found for $\bm{s}_{1}=p$ and $\bm{s}_{2}=T_{s}$.
The control performances of the optimal parametrization, $\bm{w}_{a}^{*}$, extracted from the learning curve presented in Figure 10(d) are studied next. The current scenario involves an initial over-pressurized tank operating 0.4 bar above the reference $\tilde{p}$. Figure 22 shows the time evolution of the real pressure $\check{p}$ and propellant mass $\check{m}_{l}$ over time, while the vented mass flow rate $\check{\dot{m}}_{c}$. Figures 22(b) and 22(c) show the control action over time and the corresponding pressure, respectively. Since the tank is initially in an over-pressurized state sufficiently far from $\tilde{p}$, the valve was regulated to open to its widest ($\dot{m}_{c}=0.63$ g/s), allowing the pressure to rapidly reduce until the target 5 bar are reached. This pressure drop is accompanied by a consumption of 87 kg over two roughly 45 hours. Once the target is reached, the active TVS regulates the pressure around this value. To do so, $\check{\dot{m}}_{c}$ fluctuates between -50% and 25% around its midpoint due to changes in the heat due to the disturbance, although such variations are not seen in the pressure signal.
Figure 22: T3 - Tracking performance of the pressure $\check{p}$, and variation of propellant mass $\check{m}_{l}$ over the real environment, starting from an over-pressure condition. Corresponding control action $\check{a}$ and vented mass flow rate $\check{\dot{m}}_{c}$. Thermodynamic actuation $\check{\dot{m_{v}}}$ and $\check{a}$ as a function of the pressure $p$. The acting policy $\pi(\check{\bm{e}};\bm{w}_{a})$ is retrieved by taking the optimal set of weights $\bm{w_{a}}$ from the training.
@@
Figure 23: T3 - Surface plot of the found policies, $\pi_{mf}$ (23(a)) and $\pi_{mb}$ (23(b)).
We inspect the two found policies in Figure 23. The two loops, model-based and model-free, converge to essentially the same parametrization. They show almost identical responses to the possible input space, with very minor discrepancies in the inclination of the hyperbolic tangent. Lastly, the robustness of the assimilation and control loops concludes this overview. The procedure consists in evaluating the cost of the optimal weights $\bm{w}_{p}$ and $\bm{w}_{a}$ over observations unseen during the training phase. To this end, we carried out one hundred simulations on random initial conditions and disturbances. For each of them, we computed the assimilation cost $\mathcal{L}_{p}$, as well as the control cost $\mathcal{L}_{a}$ using the model-based and model-free policies. This investigation, summarized in the violin plots presented in Figures 24(a) and 24(b), proves that the weights obtained at the end of the learning phase are valid and robust even in unseen scenarios. Moreover, in contrast to what was observed in the wind turbine test case, the model-based and model-free weights are nearly identical, thus their performances are also identical as supported by Table 2, which shows a difference between the two control performances in the order of $\mathcal{O}(10^{-2})$. This is likely a consequence of both approaches converging in the same region of the search space.
Figure 24: T3 - Robustness analysis of of model-free and model-based policies (a) and assimilated closure (b), computed over 100 episodes.
@@ 7 Conclusions and Outlooks
The performance of the Reinforcement Twinning framework, both in training the digital twin and in the training of the controlling agent, proves the framework’s potential beyond any specific application. The merit stems from the complementary characteristics of the two approaches: the model-based is extremely sample-efficient but prone to fall stuck into local minima, while the model-free is sample-inefficient but able to better explore the parameter space and avoid model limitations.
The paths towards future extensions are many. On the one hand, we could introduce a more advanced learning feedback mechanism between the model-free and the model-based methods, which is currently limited to a “cloning” procedure. This could allow, for example, different policy architecture between the two, unleashing the potential of reinforcement learning to handle arbitrarily complex policies in the form of artificial neural networks. On the other hand, we could introduce more sophisticated methods of integrating the digital twin in the training of the model-free loop, for instance, by using the virtual environment to boost the training of the critic network with strategies akin to traditional model-based reinforcement learning.